OpenStudy (anonymous):

Use the Rational Zeros Theorem to write a list of all possible rational zeros of the function.

OpenStudy (anonymous):

f(x) = -2x^4 + 4x^3 + 3x^2 + 18

OpenStudy (anonymous):

FACTOR -2x^4 + 4x^3 + 3x^2 + 18

OpenStudy (shubhamsrg):

otherwise, product of roots of this function = -18/-2 = 9 all possible factors of 9 = -9,-3,-1,1,3,9 try each of em to see if they are 0's of the function..

OpenStudy (anonymous):

All plus/minus factor combos of 18/2 = 1/1,-1/1,1/2,-1/2, 2/1, etc

OpenStudy (shubhamsrg):

ohh yes,,am sorry,,,i missed that..estudier is right..